Suppose the short run production function can be represented by

Q=60,000L^2-1000L^3 then, determine

A) the level labor employement that maximazes the level of output

B) the level of employment that maximazes APL and the maximum APL



1
Expert's answer
2022-01-02T18:20:46-0500

A) The level labor employement that maximazes the level of output is:

"MPL = Q'(L) = 120,000L - 3,000L^2 = 0,"

L(40 - L) = 0,

L = 0 (does not apply), L = 40 units.


B) The level of employment that maximazes APL is:

"APL = Q\/L = 60,000L - 1,000L^2 = 0,"

L(60 - L) = 0,

L = 0 (does not apply), L = 60 units.

The maximum APL is:

"APL = 60,000\u00d760 - 1,000\u00d760^2 = 0."